# A generalized statistical model for fits to parton distributions

### v1

@@ Abstract
Parton distribution functions (PDFs) form an essential part of particle physics calculations. Currently, the most precise predictions for these non-perturbative functions are generated through fits to global data. A problem that several PDF fitting groups encounter is the presence of tension in data sets that appear to pull the fits in different directions. In other words, the best fit depends on the choice of data set. Several methods to capture the uncertainty in PDFs in presence of seemingly inconsistent fits have been proposed and are currently in use. These methods are important to ensure that uncertainty in PDFs are not underestimated. Here we propose a novel method for estimating the uncertainty by introducing a generalized statistical model inspired by unsupervised machine learning techniques, namely the Gaussian Mixture Model (GMM). Using a toy model of PDFs, we demonstrate how the GMM can be used to faithfully reconstruct the likelihood associated with PDF fits, which can in turn be used to accurately determine the uncertainty on PDFs, especially in presence of tension in the fitted data sets. We further show how this statistical model reduces to the usual chi-squared likelihood function for a consistent data set and provide measures to optimize the number of Gaussians in the GMM.
@@ I Introduction
Parton distribution functions (PDFs) are non-perturbative functions quantifying the probability of finding a quark or gluon with certain momentum fraction $x$ from a nucleon. Taking advantage of a wide variety of experimental data and advanced computating packages for perturbative QCD (and electroweak) calculations, various groups [1, 2, 3, 4, 5] provide a sophisticated determination of PDFs. The recently released NNPDF [6] and xFitter [7] fitting codes open opportunities for general users to also conduct global QCD analyses.
@@
Many issues could be encountered when trying to perform global fits to data sets from a variety of experiments. One of these issues, which is the primary focus of this paper, is in dealing with data that appear to be in tension with each other, leading to inconsistencies within the global fit. For example, suppose we have two data sets that pull the fit parameters to different values. In this scenario, the data are in tension with each other. The source of such tension could arise due to deficiencies in theoretical models and calculations or even from underestimation of experimental uncertainties. To keep the discussion simple, we work with an example of the latter, although, the method presented here is agnostic to the source of tension.
@@
Modern PDFs perform fits to data by defining a likelihood or loss function. For PDFs produced by CTEQ-TEA (CT) [1] and MSHT [2], the likelihood is the $\Delta\chi^{2}$. Under normal circumstances with a consistent set, the uncertainty can be defined by the $68\%$ confidence level limit of the $\chi^{2}$ distribution, which amounts to varying the fit and determining boundaries where $\Delta\chi^{2}=1$. The advanced methodology adopted by both CTEQ-TEA and MSHT is to introduce a tolerance $T$ such that the fits should be varied within $\Delta\chi^{2}=T^{2}$. Values of $T$ are chosen either dynamically [2] or set to a fixed value [1]. Typical values of $T$ are much larger than $1$, which increases the uncertainties in the PDFs. NNPDF uses the Monte-Carlo method for determining uncertainties, where they produce representative sample of PDFs by sampling their likelihood or loss function with the help of pseudo-data,which might be one of the reasons for the uncertainties from NNPDF’s analysis being notably different [17]. Although, as pointed out in Refs. [18, 19, 20], the level of inconsistency with data sets is found to be negligible in the NNPDF3.1 analyses, further study is needed to understand differences between NNPDF and the other fitting groups. One such task was performed in Ref. [21], which details benchmark comparisons of three global PDF sets - CT18, MSHT20 and NNPDF3.1 - and their similarities and differences that have been observed. The end result of that study was a new PDF4LHC21 ensemble of combined PDFs suitable for a wide range of LHC applications.
In this paper we describe how to apply a more general statistical framework to accommodate potential inconsistencies in PDF data. Specifically, we show how the Gaussian Mixture Model (GMM), which is commonly used as an unsupervised machine learning technique, can be adapted to produce PDF uncertainties that faithfully represent the likelihood of PDF fits. Our major findings are summarized below for the benefit of the reader.
  • The Least Squares(LS) method is based on the assumption that the likelihood function looks like a single Gaussian distribution. Thus the likelihood is not faithfully represented when dealing with inconsistent data sets. The resulting uncertainties can be highly overestimated in some regions whereas they can simultaneously be highly underestimated in other regions.
  • The GMM provides a way to faithfully represent the likelihood of a combination of data sets with tension, which in turn gives a more accurate estimate of uncertainties. Therefore, the implementation of GMM in the PDF global fitting can be expected to accommodate the data sets with varying degrees of tension.
  • It is necessary to assess the level of tension in the data sets and to determine the optimal hyperparameters of the GMM. We use the Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) to make such a determination and avoid either over-fitting or under-fitting.
@@
The rest of the paper is organized as follows. In Sec. II, we set up a toy PDF model and describe different pseudo-data sets, some of which have obvious tension between them. Section III introduces the Gaussian Mixture Model as the statistical model to be used in order to deal with inconsistencies in data sets. In Sec. IV, the results of the GMM fit are compared with those of conventional fits using the usual least squares fit. Consistency between the GMM and the commonly-used Least Squares (LS) method is evaluated using a data set with no tension in Sec. V. Finally, we discuss our results and conclude in Sec. VI.
@@ II Tension in a toy model
In this section we introduce a toy model of PDFs and perform LS fits in order to illustrate issues that arise when there is tension between data sets.
II.1 Two pseudo-data sets with tension
We begin by utilizing the toy model of pseudo-PDF suggested in Ref. [22] and choose the functional form of the underlying truth to be,
@@ II.1 Two pseudo-data sets with tension
Here $\left\{a_{0},a_{1},a_{2},a_{3},a_{4},a_{5}\right\}$ are parameters of the function. To keep the discussion clear we will only perform fits to subsets of these parameters, while keeping others fixed. We would like to emphasize that although we choose an explicit functional form, the starting point could have also been a deep neural network or a polynomial parametrization. [^1] The use of a deep neural network is supported by the Universal Approximation Theorem [23] whereas a polynomial is supported by the Stone-Weierstrass Theorem [24, 25]. These theorems ensure that any smooth function can be well approximated either by neural networks or by polynomials. Here, we restrict ourselves to use only the functional form in Eq. (1) for clarity of presentation.
We generate pseudo-data sets by setting values of the parameters $a_{i}$ to certain fixed values. Then, for each fixed value of $x$ we generate a pseudo-data point by assuming the data follows a normal distribution about the truth with uncertainty given by
$\Delta g(x)=\frac{\alpha}{\sqrt{g(x)}}.$ (2)
$g_{D}(x)=\Big{(}1+r\times\Delta g(x)\Big{)}g(x),$ (3)
For the following analysis, we create six pseudo-data sets that are summarized later in Table 4. In this section we describe only two of those pseudo-data sets, which have different underlying truth values and thus are in tension with each other. We refer to these two data sets as pseudo-data #$\#1$ and pseudo-data #$\#2$. These pseudo-data sets are generated on the same array of $x$ values but with different values of parameters $a_{i}$ as shown in Table 1. The comparison of these two pseudo-data sets, along with that of their respective pseudo-PDF truths, is shown in Fig. 1. The tension between the pseudo-data sets is strong in the range $0.110^{-2}<x<$ 0.1, and less significant in other ranges of $x$. Generating the pseudo-data sets for the same values of $x$ ensures that the tension is explicit without any wiggle-room to smooth over tension with a flexible parameterization of pseudo-PDFs or due to a different theoretical model. It is important to note that, in general, inconsistencies in data might be a result of an inadequate theoretical model, such as underestimation of theoretical uncertainties or a less flexible parametrization of non-perturbative PDFs, or from an under-estimation of the experimental data error. In either case, the GMM method we propose here, will be able to capture uncertainties from both theoretical and experimental. Hence, we do not concern ourselves with the sources of data tension, but rather, we investigate methods to estimate uncertainty when fitting inconsistent data with a fixed theoretical prescription.
$N_{\text{pt}}$ $a_{0}$ $a_{1}$ $a_{2}$ $a_{3}$ $a_{4}$ $a_{5}$
pseudo-data #1 50 30 0.5 2.4 4.3 2.4 -3.0
pseudo-data #2 50 30 0.5 2.4 4.3 2.6 -2.8
Table 1: The choice of parameters ($a_{i}$) of the function $g(x)$ defined in Eq.(1) used to generate the two pseudo-data sets #1 and #2. $N_{pt}$ refers to the number of data points in the set.
Figure 1: Two pseudo-data sets, $g_{D}(x;\#1)$ (black) and $g_{D}(x;\#2)$ (red) are in tension with each other. The underlying truth $g(x)$ for individual pseudo-data sets are also shown by the solid curves which are referred to as truth $g(x;\#1)$ and truth $g(x;\#2)$. The plot on the left shows the pseudo-PDFs whereas the plot on the right shows their ratio with respect to the underlying truth $g(x;\#1)$.
II.2 Least Squares Fits and Tension
Before introducing the GMM method we first illustrate the method of dealing with inconsistent data in the Least Squares fitting framework. Fits are performed by optimizing the log likelihood, which in our case is the $\chi^{2}$ function defined as
@@ II.2 Least Squares Fits and Tension
To illustrate the problem with the LS fit when tension exists, three parallel LS fits are performed, which we refer to as SG-A (performed only with pseudo-data #1), SG-B (performed only with pseudo-data #2) and SG-C (performed by combining both pseudo-data #1 and pseudo-data #2). We label these fits as “SG” to refer to the single-Gaussian nature of the likelihood in LS method, in contrast to the multi-modal distribution that we will introduce later. Results of these fits are shown in Table 2. In these fits, we use the true functional form in Eq. (1), for the theory calculation, and fix parameters $\{a_{0},a_{1},a_{2},a_{3}\}$ to their true values which can be found in Table 1. The parameters $a_{4}$ and $a_{5}$ are varied and the values that minimize the $\chi^{2}$ function are shown in Table 2. The values of $\chi^{2}_{\#1}/N_{\text{pt}}$, $\chi^{2}_{\#2}/N_{\text{pt}}$, and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}$ indicate the goodness of fit to individual pseudo-data sets # 1, # 2 and the combined pseudo-data set respectively. A good fit will have values of a $\chi^{2}/N_{\text{pt}}\sim 1$, whereas a poor fit has $\chi^{2}/N_{\text{pt}}$ values significantly larger than 1. The best-fit of SG-A, where only pseudo-data set # 1 is included, is found at the minimum of $\chi^{2}_{\#1}/N_{\text{pt}}=0.88$, with values of $a_{4}=2.32$ and $a_{5}=-3.22$. Also shown in the same row for SG-A are values of $\chi^{2}_{\#2}/N_{\text{pt}}$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}$, these indicate the values of each of the likelihood functions $\chi^{2}_{\#2}/N_{\text{pt}}$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}$ determined at the best fit values of $a_{4}$ and $a_{5}$ for fits SG-A. We see that for SG-A the values of $\chi^{2}_{\#2}/N_{\text{pt}}=6.55$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}=3.72$ are significantly larger than one, indicating that there is significant tension between both the combined data set and the data set #2.
@@
Likewise, the best-fit of SG-B, where only pseudo-data set # 2 is included, is found at the minimum of $\chi^{2}_{\#2}/N_{\text{pt}}=1.02$, indicating a reasonably good fit. For SG-B, the values $\chi^{2}_{\#1}/N_{\text{pt}}$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}$ are determined by using the best fit value of $a_{4}=2.63$ and $a_{5}=2.72$. Once again large values of $\chi^{2}_{\#1}/N_{\text{pt}}=7$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}=4.01$ indicate tension with both the combined data set and data set #1. Recall, that tension between the two data-sets, pseudo-data #1 and pseudo-data #2 are generated by starting with different choices of $a_{4}$ and $a_{5}$ as given in Table 1. It can be seen, as expected, that the best-fit values of $a_{4}$ and $a_{5}$ in the two fits without tension, namely SG-A and SG-B, are indeed close to their underlying truth values.
@@
The fit, SG-C, to the the combined data set where both pseudo-data sets # 1 and # 2 are included, is found at the minimum of $\chi^{2}_{\#1+\#2}/N_{\text{pt}}=2.42$, indicating, as expected, a poor fit to the combined data set. In this case, the values of $\chi^{2}_{\#1}/N_{\text{pt}}=2.27$, $\chi^{2}_{\#2}/N_{\text{pt}}=2.42$ indicate that the combined fit, still has tension with both individual data sets.
@@
Although the fit to the combined data is poor, the main problem arises when attempting to determine the relevant uncertainty associated with such a fit. This is illustrated in Fig. 2, which presents Results of the LS fits SG-A, SG-B and SG-C. For all the three fits, results are presented in the form of replica sets, where the probability distribution of the pseudo-PDF is sampled with the Monte Carlo method as described in [6]. For the purpose of comparison, we also estimate the uncertainty of SG-C fit using the Hessian method [1], which is consistent with the Monte Carlo method. Here we only consider the Hessian uncertainty generated by the variation of $\chi^{2}$ and determine the uncertainty with the criterion that $\Delta\chi^{2}=1$. Here, we do not include any additional tolerance criterion, such as the two-tier penalty for the CT uncertainty [1] and the dynamic tolerance for the MSHT uncertainty [2]. Later, we will include such tolerance criteria for comparison with the GMM.
@@
In panels 2 and 2, pseudo-PDFs from fits SG-A, SG-B and SG-C are found to be in disagreement in the range $0.210^{-3}<x<$ 0.2. Even the $1\sigma$ uncertainty bands do not overlap in this region. For $x>0.2$, the tension in data is weak due to larger uncertainty in the data and this can be seen from the overlap of the uncertainty bands in this region. Even though the uncertainty bands overlap at large $x$, the impact of significant tension in the range $0.210^{-3}<x<$ 0.2 still affects pseudo-PDF at large $x$ as can be observed by the different shapes of the uncertainty bands for the different LS fits. For the LS fit SG-C, the Hessian uncertainty with $\Delta\chi^{2}=1$ is also provided, which deviates from the Monte Carlo uncertainty by a negligible amount. The tension between the different fits, SG-A, SG-B and SG-C, can also be observed by looking at the result of the fits to the parameters $a_{4}$ and $a_{5}$, shown in Fig. 2. As expected, the true values of the parameters $a_{4}$ and $a_{5}$ fall within the regions covered by the Monte Carlo sample of 1000 replicas for each of the fits SG-A and SG-B. However, the true values of the fit parameters are far from any of the replicas generated with SG-C. Furthermore, none of the replicas generated for either of the fits overlap and to make matters worse, neither the $1\sigma$ nor $3\sigma$ uncertainty bands of SG-C, generated using the Hessian method, overlap with the two truth values for $a_{4}$ and $a_{5}$ shown in Table 1. The tolerance criteria used by the PDF fitting groups would increase the uncertainty, but as to be shown in Section IV, this approach will not be sufficient to incorporate both data sets into a single fit.
One might wonder why the fit SG-C, which combines both data sets, is unable to reproduce the desired uncertainty bands that should look like some combination of the uncertainty bands of both SG-A and SG-B. Such a combination of uncertainty bands would arise only if the likelihood function were multi-modal. On the other hand, fit SG-C assumes that there is only a single mode for the likelihood. In other words, we have ignored the two individual likelihoods generated by the two data sets and tried to define a new likelihood in SG-C that we have assumed to be uni-modal. Another way of stating this is to say that a naive combination of Monte-Carlo samples from SG-A and SG-B, seen in Fig. 2, is a better representation of the combined likelihood of both data sets. This combined likelihood is multi-modal, unlike that of SG-C. Of course, just by adding the Monte Carlo samples does not suffice and we need a proper method to combine the samples. In order to combine likelihoods in a statistically robust way, we introduce the GMM method in the next section which provides a way to combine likelihoods into a multi-modal distribution.
Before we conclude this section, we comment on alternate methods to deal with such inconsistencies in fits. For example, the Particle Data Group (PDG) [26], recommends increasing the uncertainty on data sets, either all data sets if they are of similar precision or the data set from the experiment with the highest precision, by a scale factor $S=\sqrt{\chi^{2}/N_{\text{pt}}}$. For data sets with tension, $\sqrt{\chi^{2}/N_{\text{pt}}}>1$ and hence the scale factor $S>1$, implying an increase in uncertainty. For a large data set ($N_{\text{pt}}\gg 1$), increasing the uncertainty on all data sets implies that $\chi^{2}/N_{\text{pt}}\simeq 1$, thus ensuring a good fit. This idea has been further explored and elucidated in both the Bayesian and Frequentist approach [27, 28, 29]. The GMM presented here, not only provides an alternate way to combine inconsistent data sets without increasing their uncertainty, but also, if desired, provides an alternate estimate for the scale factor $S$ required to make data sets consistent. We describe some of these ideas briefly in Sec. V.2.

### v2

@@ Abstract
Parton distribution functions (PDFs) form an essential part of particle physics calculations. Currently, the most precise predictions for these non-perturbative functions are generated through fits to global data. A problem that several PDF fitting groups encounter is the presence of tension in data sets that appear to pull the fits in different directions. In other words, the best fit depends on the choice of data set. Several methods to capture the uncertainty in PDFs in presence of seemingly inconsistent fits have been proposed and are currently in use. These methods are important to ensure that uncertainty in PDFs are not underestimated. Here we propose a novel method for estimating the uncertainty by introducing a generalized statistical model based on Bayesian Hierarchical models which is implemented via the Gaussian Mixture Model (GMM). The methodology is inspired by unsupervised machine learning techniques and is closely related to the statistical methods of Ensemble Learning and Bayesian Model Averaging. Using a toy model of PDFs, we demonstrate how the GMM can be used to faithfully reconstruct the likelihood associated with PDF fits, which can in turn be used to accurately determine the uncertainty on PDFs, especially in the presence of tension in the fitted data sets. We further show how this statistical model reduces to the usual chi-squared likelihood function for a consistent data set and provide measures to optimize the number of Gaussians in the GMM.
@@ I Introduction
Parton distribution functions (PDFs) are non-perturbative functions quantifying the probability of finding a quark or gluon with certain momentum fraction $x$ from a nucleon. Taking advantage of a wide variety of experimental data and advanced computing packages for perturbative QCD (and electroweak) calculations, various groups [1, 2, 3, 4, 5] provide a sophisticated determination of PDFs. The recently released NNPDF [6] and xFitter [7] fitting codes open opportunities for general users to also conduct global QCD analyses.
@@
Many issues could be encountered when trying to perform global fits to data sets from a variety of experiments. One of these issues, which is the primary focus of this paper, is in dealing with data that appear to be in tension with each other, leading to inconsistencies within the global fit. For example, suppose we have two data sets that pull the fit parameters to different values. In this scenario, the data are in tension with each other. The source of such tension could arise due to deficiencies in theoretical models and calculations or even from incorrect estimation of experimental uncertainties. To keep the discussion simple, we work with an example of the latter, although, the method presented here is agnostic to the source of tension.
@@
Modern PDFs perform fits to data by defining a likelihood or loss function. For PDFs produced by CTEQ-TEA (CT) [1] and MSHT [2], the likelihood is the usual $\Delta\chi^{2}$. Under normal circumstances with a consistent set, the uncertainty can be defined by the $68\%$ confidence level limit of the $\chi^{2}$ distribution, which amounts to varying the fit and determining boundaries where $\Delta\chi^{2}=1$.
The convention used by the PDF global analysis groups, such as CT and MSHT, is to introduce a tolerance $T$ such that the fits should be varied within $\Delta\chi^{2}=T^{2}$. Values of $T$ are chosen either dynamically [2] or set to a fixed value [1]. Typical values of $T$ are much larger than $1$, which increases the uncertainties in the PDFs. For example, $T^{2}=37$ corresponds to the CT tolerance at the 68% CL.
@@
NNPDF uses the Monte-Carlo method for determining uncertainties, where they produce representative sample of PDFs by sampling their likelihood or loss function with the help of pseudo-data,which might be one of the reasons for the uncertainties from NNPDF’s analysis being notably different [17]. Although, as pointed out in Refs. [18, 19, 20], the level of inconsistency with data sets is found to be negligible in the NNPDF3.1 analyses, further study is needed to understand differences between NNPDF and the other fitting groups. One such task was performed in Ref. [21], which details benchmark comparisons of three global PDF sets - CT18, MSHT20 and NNPDF3.1 - and their similarities and differences that have been observed. The end result of that study was a new PDF4LHC21 ensemble of combined PDFs suitable for a wide range of LHC applications.
In this paper we describe how to apply a more general statistical framework, namely that of Bayesian Hierarchical Models [22, 23, 24], to accommodate potential inconsistencies in PDF data. This rather non-trivial task is accomplished with the aid of the Gaussian Mixture Model (GMM), which is commonly used as an unsupervised machine learning technique as well as a method of combining models. The approach used here can be interpreted as an implementation of the statistical method of Ensemble Learning [25, 26] and the related method of Bayesian Model Averaging (BMA) [27, 28, 29, 30, 31]. BMA is a statistical method used to account for uncertainties arising from model selection. Although it is most often used in the context of linear regression, here we apply it to highly non-linear fits. [^1] Application of Bayesian Model Averaging to analyze Lattice Field Theory results can be found in Refs. [32, 33, 34]. On the other hand, Ensemble Learning, also known as Ensemble Averaging or Bayesian Model Combination, is a method to combine predictions of multiple statistical models. Although there are some subtle differences between Ensemble Learning and Bayesian Model Averaging, which we discuss briefly in Appendix B.2, we do not dwell on these differences as they are subject to differences in the interpretation of priors. Ensemble Learning is primarily used to combine different theoretical models and their point estimates. On the other hand, Bayesian Hierarchical Models are primarily applied to cases where there are variations across groups of data and its use has also been suggested in estimating uncertainty when there is tension in data sets [24]. Here, we demonstrate how Ensemble Learning can be used to construct a Bayesian Hierarchical Model and how it can be used to provide an estimate of the uncertainty in the scenario where there are tensions among data sets. We will not discuss combining or averaging over different theoretical models which is beyond the scope of this paper.
Our major findings are summarized below for the benefit of the reader.
  • When fitting inconsistent data sets with tension, the usual Least Squares (LS) method, that employs the $\chi^{2}$ likelihood, underestimates uncertainty. Specifically, the posterior probability distribution for fitted theory parameters should reflect the tension in the data sets and appropriately span the range of possible theory values. We demonstrate how the LS method and its adaptations to address tension in data sets do not appropriately span the range of possible theory values. Another drawback of these methods is that they introduce ad hoc criteria to expand uncertainties. These deficiencies are overcome by the use of the GMM.
  • We introduce a new method to fitting inconsistent data, utilizing the Gaussian Mixture Model. The method is closely related to Ensemble Learning and Bayesian Model Averaging, that are used to estimate uncertainties due to selection of theoretical models. Here, we show how these methods can be repurposed to help determine uncertainties by constructing a Bayesian Hierarchical Model when there is tension between data sets.
  • The main idea is a modification of the likelihood function used to determine the posterior probability distribution of the theory parameters. This modification, given in Eq. (III), accounts for the possibility of fitting with a multi-modal distribution rather than a uni-modal distribution assumed in the LS method. Estimation of theory parameters and their associated uncertainties proceed through the usual Maximum Likelihood estimation. We describe methods and provide formulae for the estimate of the mean and variance in Eq. (12) and Eq. (20), respectively. We demonstrate the use and characteristics of the GMM method using various toy examples where we try to combine various pseudo-data sets with varying degrees of tension.
  • The fit presented in Fig. 4 and Fig. 4 is an example of how the GMM method provides a more accurate representation of the uncertainty compared to the LS method by spanning the range of possible theory predictions, given the tension in the data sets. Specifically, using the GMM, we find that the uncertainty bands are larger when the tension is more prominent and smaller otherwise. On the other hand, more traditional methods of addressing tension tend to expand uncertainties more uniformly, resulting in larger uncertainty bands in regions where the tension is not large.
  • The GMM introduces the number of modes ($K$) of the distribution as a hyperparameter. We demonstrate the use of Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) not only to determine $K$ in order to avoid over-fitting or under-fitting, but also to be used as quantitative tests of the consistency of data sets.
@@
The rest of the paper is organized as follows. In Sec. II, we set up a toy PDF model and describe different pseudo-data sets, some of which have obvious tension between them. In Appendix A we describe how the pseudo-data sets we use in the fits are generated. Section III introduces the Gaussian Mixture Model as a statistical model that can be used to tackle inconsistencies in data sets. In Sec. IV, the results of the GMM fit are compared with those of conventional fits using the usual least squares fit. Consistency between the GMM and the commonly-used Least Squares (LS) method is evaluated using a data set with no tension in Sec. V. We discuss our results and conclude in Sec. VI. Additionally, we describe a more intuitive Bayesian perspective of the GMM in Appendix B and discuss some of the differences and intricacies related to BMA and Ensemble Learning in Appendix B.3. In Appendix C, we provide a much simpler and more accessible example using the real world data of $W$-boson mass measurements.
@@ II Tension in a toy model
In this section, we introduce a toy model of PDFs and the pseudo-data sets to fit the PDFs. We perform LS fits to illustrate the problems encountered when there is tension between data sets.
Figure 1: Two pseudo-data sets, Pseudo-Data #1 (black) and Pseudo-Data #2 (red) are shown to be in tension with each other. The solid curves which are referred to as g(x;#1)g(x;\#1) and g(x;#2)g(x;\#2), correspond to the curves given by the function in Eq. (1), with parameter values given in Table 4. The plot on the left shows the pseudo-PDFs whereas the plot on the right shows their ratio with respect to g(x;#1)g(x;\#1).
We begin by considering two sets of data that we refer to as pseudo-data #1 and pseudo-data #2, containing fifty data points each. These two data sets are not real, but are simulated. However, we treat them as though they are measurements that likely came from two separate experiments. Details of how these pseudo-data sets simulated are relegated to Appendix A. The two pseudo-data sets are presented in Fig. 1, showing measured values of a pseudo-PDF as a function of $x$. Both data sets have a larger uncertainty at low and high values of $x$, but smaller uncertainties at intermediate values of $x\sim 10^{-2}$, to mimic general features of real experimental data. This choice represents the possibility that larger values of the PDF may correspond to larger number of events and therefore a reduced statistical uncertainty. The data sets have an obvious tension between them in the region $10^{-2}\lesssim x\lesssim 10^{-1}$. Such a tension might be due to underestimated systematics or even due to underestimated theoretical uncertainties.
Given the two pseudo-data sets, we would like to perform a combined fit to both data sets. We fit the data to a function of the form
@@
Here $\left\{a_{0},a_{1},a_{2},a_{3},a_{4},a_{5}\right\}$ are the theory parameters that we would like to determine from the pseudo-data. In addition, we would also like to estimate the uncertainty on these fitted parameters from the given data. To keep the discussion simple we only perform fits to a subset of these parameters, while keeping others fixed. We would like to emphasize that although we choose an explicit functional form, the starting point could have also been a deep neural network or a polynomial parametrization. [^2] The use of a deep neural network is supported by the Universal Approximation Theorem [35] whereas a polynomial is supported by the Stone-Weierstrass Theorem [36, 37]. These theorems ensure that any smooth function can be well approximated either by neural networks or by polynomials. Here, we restrict ourselves to use only the functional form in Eq. (1) for clarity of presentation.
Our purpose is to determine point estimates, i.e. the mean and variance, of the theoretical parameters $a$, given the experimentally measured data. Note, in the process of generating the two pseudo-data sets, we do have knowledge the values of these parameters, c.f. Appendix A. However, when performing the fits we do so by excluding any prior knowledge of the theory parameters $a$. Our purpose is not to try to recover parameter values that were used to generate each pseudo-data set separately, instead, we would like to determine an estimate of the theory parameters $a$ given a combination of both data sets. Furthermore, to be consistent with QCD, we demand that we have only one set of theoretical parameters, so that there is only one estimate of the mean. We now proceed to first perform LS fits to the two pseudo-data sets in Fig. 1.
II.1 Least Squares Fits and Tension
Before introducing the GMM method we first illustrate the method of handling inconsistent data in the Least Squares fitting framework. Fits are performed by optimizing the log likelihood, which in our case is the $\chi^{2}$ function defined as
@@ II.1 Least Squares Fits and Tension
To illustrate the problem with the LS fit when tension exists, three LS fits are performed, which we refer to as SG-A (performed only with pseudo-data #1), SG-B (performed only with pseudo-data #2) and SG-C (performed by combining both pseudo-data #1 and pseudo-data #2). We label these fits as “SG” to refer to the single-Gaussian nature of the likelihood in the LS method, in contrast to the multi-modal distribution that we will introduce later. Results of these fits are shown in Table 1.
We fit the data to the function given in Eq. (1), by varying and determining the parameters of the function that minimize the $\chi^{2}$ in Eq. (2). To keep the discussion and fits simple in this toy example, we fix parameters $\{a_{0},a_{1},a_{2},a_{3}\}$ to values which can be found in Table 4. Only two parameters $a_{4}$ and $a_{5}$ are varied and the values that minimize the $\chi^{2}$ function are shown in Table 1. The values of $\chi^{2}_{\#1}/N_{\text{pt}}$, $\chi^{2}_{\#2}/N_{\text{pt}}$, and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}$, which are approximately equal to the reduced $\chi^{2}$, indicate the goodness of fit to individual pseudo-data sets # 1, # 2 and the combined pseudo-data set, respectively. As is well known, a good fit will have values of a $\chi^{2}/N_{\text{pt}}\sim 1$, whereas a poor fit has $\chi^{2}/N_{\text{pt}}$ values significantly larger than 1.
The best-fit of SG-A, where only pseudo-data set # 1 is included, is found with a minimum of $\chi^{2}_{\#1}/N_{\text{pt}}=0.88$, with values of $a_{4}=2.32$ and $a_{5}=-3.22$. Also shown in the same row of Table 1 for SG-A are values of $\chi^{2}_{\#2}/N_{\text{pt}}$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}$. These indicate the values of each of the likelihood functions $\chi^{2}_{\#2}/N_{\text{pt}}$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}$ determined at the best fit values of $a_{4}$ and $a_{5}$ for fits SG-A. We see that for SG-A the values of $\chi^{2}_{\#2}/N_{\text{pt}}=6.55$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}=3.72$ are significantly larger than one, indicating that there is significant tension between both the combined data set and data set #2.
@@
Likewise, the best-fit of SG-B, where only pseudo-data set # 2 is included, is found with a minimum $\chi^{2}_{\#2}/N_{\text{pt}}=1.02$, indicating a reasonably good fit. For SG-B, the values $\chi^{2}_{\#1}/N_{\text{pt}}$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}$ are determined by using the best fit value of $a_{4}=2.63$ and $a_{5}=2.72$. Once again large values of $\chi^{2}_{\#1}/N_{\text{pt}}=7$ and $\chi^{2}_{\#1+\#2}/N_{\text{pt}}=4.01$ indicate tension with both the combined data set and data set #1.333 The two pseudo-data sets, #1 and #2, are generated by starting with different choices of $a_{4}$ and $a_{5}$ as given in Table 4. As expected, these values of $a_{4}$ and $a_{5}$ are similar to the fitted values in Table 1 for SG-A and SG-B and lie within the 1-$\sigma$ confidence interval of the fit.
@@
The fit, SG-C, to the the combined data set where both pseudo-data sets # 1 and # 2 are included, is found with a minimum $\chi^{2}_{\#1+\#2}/N_{\text{pt}}=2.42$, indicating, as expected, a poor fit to the combined data set. In this case, the values of $\chi^{2}_{\#1}/N_{\text{pt}}=2.27$, $\chi^{2}_{\#2}/N_{\text{pt}}=2.42$ indicate that the combined fit, still has tension with both individual data sets.
@@
Although the fit to the combined data is poor, the main problem arises when attempting to determine the relevant uncertainty associated with such a fit. This is illustrated in Fig. 2, which presents results of the LS fits SG-A, SG-B and SG-C. For all three fits, results are presented in the form of replica sets, where the probability distribution of the pseudo-PDF is sampled with the Monte Carlo method as described in [6]. For the purpose of comparison, we also estimate the uncertainty of the SG-C fit using the Hessian method [1], which is consistent with the Monte Carlo method. Here we only consider the Hessian uncertainty generated by the variation of $\chi^{2}$ and determine the uncertainty with the criterion that $\Delta\chi^{2}=1$. We do not include any additional tolerance criterion, such as the two-tier penalty for the CT uncertainty [1] and the dynamic tolerance for the MSHT uncertainty [2]. Later, we will include such tolerance criteria for comparison with the GMM.
@@
In panels 2 and 2, pseudo-PDFs from fits SG-A, SG-B and SG-C are found to be in disagreement in the range $10^{-3}<x<0.2$. Even the $1\sigma$ uncertainty bands do not overlap in this region. For $x>0.2$, the tension is weak due to larger uncertainty in the data which is represented by the overlapping uncertainty bands in this region. Even though the uncertainty bands overlap at large $x$, the impact of significant tension in the range $10^{-3}<x<0.2$ still affects the pseudo-PDF at large $x$ as can be observed by the different shapes of the uncertainty bands for the different LS fits. For the combined LS fit SG-C, the Hessian uncertainty with $\Delta\chi^{2}=1$ is also provided, which deviates from the Monte Carlo uncertainty by a negligible amount. The tension between the different fits, SG-A, SG-B and SG-C can also be observed by looking at Fig. 2 that shows uncertainty on the parameters $a_{4}$ and $a_{5}$. Monte Carlo samples of 1000 replicas are generated for each of the fits SG-A, SG-B and SG-C.
None of the replicas generated for either of the fits overlap, once again indicating the strong tension among the psuedo-data sets.
We would like to emphasize that either or both experiments may have underestimated systematic uncertainties and in the absence of further information about the source of systematics, the range of possible theory values should span the theory values predicted by each experimental data set individually. For example, an uncertainty band that looks like some combination of the uncertainty bands of both SG-A and SG-B, should envelope the central values of the fits SG-A and SG-B, or at least come close to doing so. Such an uncertainty band would be more representative of the tension in the data. Fig. 2 shows that this is not the case for the SG-C fit. The tolerance criteria used by the PDF fitting groups increases the uncertainty, but as will be shown in Section IV, this approach is not as precise as using the GMM. Further, in Sec. III, we implement Ensemble Learning with the use of the GMM and show how this allows for the combined likelihood of the pseudo-data to be multi-modal and results in a uncertainty band that overlaps with the uncertainty bands of both SG-A and SG-B.